OpenStudy (anonymous):

Help! I give medals:) Pic below The volume of two similar cylinders are 2187pi in cubed squared and 8000pi cubed in squared find...

OpenStudy (primeralph):

I'm back. Bring it on!

OpenStudy (anonymous):

 \[\frac{ Area I }{ Area II }=\frac{ a ^{2} }{ b ^{2} }\] \[\frac{ r1 }{ r2 }=\frac{ h1 }{ h2 }=\frac{ a }{ b}\] \[\frac{ Volume I }{ Volume II }=\frac{ a ^{3} }{ b ^{3} }\]